OpenStudy (anonymous):

A roller-coaster car starts at a height of 90 m at Point A and is moving at a speed of 20 m/s on a later hill at Point C. Assuming friction does not affect the motion of the car, what is its height at Point C?

OpenStudy (anonymous):

Here is the first one @wolfe8 :)

OpenStudy (wolfe8):

So energy is conserved. So at C, it will have kinetic energy and potential energy totaling to the potential energy at A. Can you set up the equations?

OpenStudy (anonymous):

I think so! One sec...

OpenStudy (anonymous):

Is this an instance where we are trying to find deltaK?

OpenStudy (anonymous):

Or is it just K = 1/2 m * v^2?

OpenStudy (wolfe8):

That is the kinetic energy. That plus the potential energy mgh at C will give you mgh at A

OpenStudy (anonymous):

OH okay, so it's 1/2 *m * v^2 + mgh = PE at A

OpenStudy (anonymous):

But they don't give me m :/

OpenStudy (wolfe8):

Yup! And PE=mgh. To make it not too confusing, you can write it this way: \[mgh _{A}=\frac{ 1 }{ 2 }mv ^{2}+mgh _{C}\]

OpenStudy (wolfe8):

And since m is present on every term in the equation, you can cancel it out :)

OpenStudy (anonymous):

OH Cool! So.. 9.8 * 90 = 1/2 * 20 ^2 + 9.8 * hc

OpenStudy (anonymous):

That was easy! Great :)! Thanks. I'll start the new one

OpenStudy (wolfe8):

Yup!

Directrix (directrix):

@Daniellelovee